A random sample of 150 people was taken from a very large population. About 100 of the people in the sample were female. The standard error of the proportion is?
Answer #1

Number of Items of Interest,x =100

Sample Size,n = 150

Sample Proportion , p̂ = x/n = 100/150 = 0.6667

Standard Error , SE = √[p̂(1-p̂)/n] = √(0.6667 * (1 - 0.6667) / 150) = 0.0385